# New technique for parameter estimation and improved fits to experimental data for a set of compound Poisson distributions

### v1

@@ 5.9 Fracker and Brischle (1944)
Fracker and Brischle (1944) fitted multiple sets of data using the Poisson distribution and the Neyman Type A distribution (which they called the ‘contagious distribution’). Let us denote Neyman Type A by ‘NTA’ for brevity. We fit five of their datasets below (Fracker and Brischle (1944) Table II: Kaniksu, Mt. Spokane, Clearwater and Table III: Beaver Creek and Cow Creek). They stated that neither Poisson nor NTA fitted their data well. They obtained the best fit by using a linear combination of the two distributions, viz. ($\frac{1}{3}\times$Poisson +$\frac{2}{3}\times$NTA). (Note: they used (0.4×0.4$\times$0.4 ×Poisson +0.6×0.6$\times$0.6 ×NTA) for the last dataset, Cow Creek.) I confirmed their fits numerically. However, I obtained much better fits using the geometric Poisson distribution. The values of the relevant parameter estimators $(\hat{\lambda},\hat{p})$ for the geometric Poisson distribution were obtained using eq. (10) (and $\hat{q}=1-\hat{p}$). The sample variance was calculated using eq. (7). My fits to the five datasets are displayed in Fig. 18.
  1. 1.
    Kaniksu: The method of moments yields $(\hat{\lambda},\hat{p})=(0.5363,0.5448)$. I obtained a better fit using $(\hat{\lambda},\hat{p})=(0.47,0.6228)$. The fit is displayed in the top panel of Fig. 18.
  2. 2.
    Mt. Spokane: The method of moments yields $(\hat{\lambda},\hat{p})=(0.8197,0.2197)$. I obtained a better fit using $(\hat{\lambda},\hat{p})=(0.62,0.4631)$. The fit is displayed in the second panel of Fig. 18.
  3. 3.
    Clearwater: The method of moments yields $(\hat{\lambda},\hat{p})=(0.7095,0.4402)$. I obtained a better fit using $(\hat{\lambda},\hat{p})=(0.63,0.5608)$. The fit is displayed in the third panel of Fig. 18.
  4. 4.
    Beaver Creek: The method of moments yields $(\hat{\lambda},\hat{p})=(0.4815,0.5496)$. I obtained a better fit using $(\hat{\lambda},\hat{p})=(0.45,0.5830)$. The fit is displayed in the fourth panel of Fig. 18.
  5. 5.
    Cow Creek: The method of moments yields $(\hat{\lambda},\hat{p})=(0.4768,1.0910)$. I obtained a better fit using $(\hat{\lambda},\hat{p})=(0.56,0.9174)$. The fit is displayed in the bottom panel of Fig. 18.
I did not employ the power spectrum in this section. My main conclusion is that the datasets by Fracker and Brischle (1944) are really fitted by the geometric Poisson distribution, not Poisson and/or Neyman Type A. One can observe visually that in all five datasets in Fig. 18, the geometric Poisson distribution fits their data almost perfectly. There is some experimental scatter in the last dataset (Cow Creek), but the fitted curve interpolates between the scatter, which is the best that can be expected.

### v2

@@ 5.9 Fracker and Brischle (1944)
Fracker and Brischle (1944) fitted multiple sets of data using the Poisson distribution and the Neyman Type A distribution (which they called the “contagious distribution”). Let us denote Neyman Type A by “NTA” for brevity. We fit five of their datasets below (Table II: Kaniksu, Mt. Spokane, Clearwater and Table III: Beaver Creek and Cow Creek). They stated that neither Poisson nor NTA fitted their data well. They obtained the best fit by using a linear combination of the two distributions, viz. ($\frac{1}{3}\times$Poisson +$\frac{2}{3}\times$NTA). (Note: for Cow Creek they used (0.4×0.4$\times$0.4 ×Poisson +0.6×0.6$\times$0.6 ×NTA).) We confirmed their fits numerically. However, we obtained much better fits using the geometric Poisson distribution. Using eq. (11) yielded better fits than the method of moments eq. (10). Our values for the estimators $(\hat{\lambda},\hat{p})$ for the datasets, calculated using eq. (11), are tabulated in Table 7. The resulting fits to the data are displayed in Fig. 18. The power spectrum was not employed in this section. The main conclusion is that the datasets of Fracker and Brischle (1944) are really fitted by the geometric Poisson distribution, not a linear combination of Poisson and Neyman Type A. One can observe visually that in all five datasets in Fig. 18, the geometric Poisson distribution fits the data almost perfectly. There is some experimental scatter in the last dataset (Cow Creek), but the fitted curve interpolates between the scatter, which is the best that can be expected.